OpenStudy (anonymous):

Create your own radical function and graph that function using technology. Can someone help me come up with one?

jimthompson5910 (jim_thompson5910):

Do you have any radical functions in mind?

OpenStudy (anonymous):

Not really.

OpenStudy (abb0t):

I think the cosine graph is pretty radical :P

jimthompson5910 (jim_thompson5910):

Any radical function has a square root in it, a cube root, a 4th root, etc, so here are some examples of radical functions \[\Large y = \sqrt{x+3}\] \[\Large y = \sqrt[3]{x^2+5}\] \[\Large y = \sqrt[5]{2x-3}\]

jimthompson5910 (jim_thompson5910):

So all you have to do is come up with something like that, then graph it using a graphing calculator

OpenStudy (anonymous):

Oh okay. thank you.

jimthompson5910 (jim_thompson5910):

np